7 A woman has available 1 A, 3A, 5A, 10 A and 13 A fuses. Explain which fuse she should use for a 120V, 450 W hairdryer.
Answer #1

P = VI

I = P/V = 450 W / 120 V = 3.75 Amp

she should use fuses of capacity more than 3.75 Amp.

so she can use 5A, 10A & 13 A fuses.